OpenStudy (anonymous):

MEDALS What is the value of the expression 2^4 – 3^3?

OpenStudy (anonymous):

Are you allowed to use a calculator 2x2x2x2 - 3x3x3 16 - 27 = -11 Does that make sense?

OpenStudy (cwrw238):

2^4 = 2*2*2*2 3^3 = 3*3*3

OpenStudy (anonymous):

no im not

OpenStudy (anonymous):

Are you allowed to use a calculator 2x2x2x2 - 3x3x3 = 16 - 27 =-11 Does that make sense?

OpenStudy (anonymous):

yesh thnx